OpenStudy (anonymous):

can someone help me with #10 and #11

OpenStudy (anonymous):

okay...AC = BD... so, 8x - 3 = 2x + 9...and then. solve for x!

OpenStudy (anonymous):

and for #11, it is a similar triangles problem...u will have 2 equations and 2 unknowns... 4x - y = 1 -x + 4 = y Then solve it!

OpenStudy (anonymous):

fo 10 do i add 8x to 2x?

OpenStudy (anonymous):

no...u subtract... i will rearrange for u.. 8x - 2x = 9 + 3 .hopefully, this will make it easier for u...

OpenStudy (anonymous):

AC = BD 8x - 2x = 9 + 3 -2x -2x 6x=12 6 6 x=2

OpenStudy (anonymous):

is this correct for 10?

OpenStudy (anonymous):

how do i solve 11?

OpenStudy (anonymous):

@cfraser007

OpenStudy (anonymous):

you are correct for 10!

OpenStudy (anonymous):

for 11, u have the equations 4x - y = 1 -x + 4 = y Then, what u can do is simply get that y of the second equation to the left and the 4 to the right and then combine the equations!

OpenStudy (anonymous):

so u will have to subtract the equations  then, i think its easy to know what x can be! after that, just put that value of "x" in any of the equations to solve for y!